show that the points (1,1) (0,-3) and (9,6) are the vertices of the square and find its area.

Answer by robertb(5830)

The slope thru (1,1) and (0,-3): (1--3)/(1-0) = 4/1 = 4.
The slope thru (1,1) and (9,6): (6-1)/(9-1) = 5/8.
The slope thru (0,-3) and (9,6): (6--3)/(9-0) = 9/9 = 1.
The points couldn't be vertices of a square: none of the segments are perpendicular. (None of them are negative reciprocals of each other!)